Sherry has a goal of retiring with $266676 by making weekly deposits into an investment account whose annual interest rate is 1%. If she will retire in 22 years, how much interest will she earn? Round your answer to the nearest dollar.
Answer #1
Calculation of interest earned over 22 years
Annual Rate 1%
Since a weekly payment is being made,
A Rate per period (Annual Rate divided by 52 weeks) 0.000192308
B Future Value (FV) $266,676
Period (P) 22 years
C Period in weeks (P*52) 1144
E Weekly deposit (Use PMT function) $208.43
Formula is '=PMT (A, C, 0, B)'
D Total deposit made in 22 years (E*C)= $238,441.90
B-D Interest Earned= $28,234.1
Rounded to nearest dollar $28,234